A school has fewer than 200 pupils , when they line up in 4s there is one extra pupil and when they line up in 5s there r two extra pupils How many pupils are ther?
Answer by stanbon(75887)

A school has fewer than 200 pupils , when they line up in 4s there is one extra pupil and when they line up in 5s there r two extra pupils How many pupils are there?
----------------
4x + 1 = 200
5y + 2 = 200
-------
4x + 1 = 5y + 2
-----
5y = 4x - 1
---
y = (4/5)x - (1/5)
-------
Find whole number values for x and y::
x = 49 ; y = 39
# of pupils = 4*49+1 = 197
-----
Note:: There are many possible answers. The one I
used is less than 200 but maximizes the number
of students.
Cheers,
Stan H.
---------------